OpenStudy (anonymous):

cos 2θ = 13 sin θ − 6 solve the equation

OpenStudy (anonymous):

We have: cos(2θ) = 1 - 2sin^2θ. So, the equation becomes: 1 - 2sin^2θ = 13sinθ - 6 ==> 2sin^2θ + 13sinθ - 7 = 0. The left side factors to (2sinθ - 1)(sinθ + 7), so: (2sinθ - 1)(sinθ + 7) = 0. By the zero-product property: 2sinθ - 1 = 0 and sinθ + 7 = 0 ==> sinθ = 1/2 and sinθ = -7. Since -1 ≤ sinθ ≤ 1 for all real θ, sinθ = -7 has no real solutions. From the unit circle, sinθ = 1/2 solves to yield θ = π/6 + 2πk and θ = 5π/6 + 2πk, where k is an integer. Therefore, the solutions are θ = π/6 + 2πk and θ = 5π/6 + 2πk, where k is an integer. Source @yahooanswers

OpenStudy (anonymous):

thank you!